# The Angular Momentum Penrose Inequality

### v2

@@ 6.4 Monotonicity
Conclusion. Combining Steps 5–7, we obtain uniform $C^{1,\alpha}$ bounds for some $\alpha>0$ (possibly small but positive), independent of $p\in(1,2]$. ∎
@@
    $C^{1,\alpha}$ regularity: $\|u_{p}\|_{C^{1,\alpha}(K)}\leq C(K)$ uniformly in $p\in(1,2]$ (Lemma 6.28);
@@
These three bounds ensure that the Tolksdorf stability estimate for $p$-harmonic functions [52, Theorem 3.2] applies with constants independent of $p$, validating the Moore–Osgood double limit interchange in Remark 6.27.
@@
(i) By Sard’s theorem applied to the $C^{1,\alpha}$ function $u_{p}$ (Tolksdorf regularity [52]), the set of critical values $\{t:\exists x\in u_{p}^{-1}(t)\text{ with }\nabla u_{p}(x)=0\}$ has measure zero in $[0,1]$. For $p$-harmonic functions, the Harnack inequality [50] implies that critical points are isolated unless $u_{p}$ is constant. Since $u_{p}$ ranges from 0 to 1, it is non-constant, so $\mathcal{Z}_{p}$ is a discrete (hence measure-zero) set.
(ii) Away from $\mathcal{Z}_{p}$, the $p$-harmonic equation is uniformly elliptic. The Harnack inequality for $p$-harmonic functions [50, Theorem 1.2] gives:
@@
AMO boundary conditions exclude critical curves. For the AMO potential $u_{p}:\tilde{M}\to[0,1]$ with $u_{p}|_{\Sigma}=0$ and $u_{p}\to 1$ at infinity, we have stronger control. The key observation is that $u_{p}$ is a capacitary potential—it minimizes the $p$-energy among functions with the given boundary values. By Manfredi [37, Theorem 4.1], capacitary potentials in dimension 3 have critical sets of dimension $\leq 0$ (isolated points) when the boundary data is “generic” in the sense that no boundary component has vanishing $p$-capacity.
@@
Saddle points of capacitary potentials are isolated by the classification of singularities in Aronsson–Lindqvist [9, Section 5]. Therefore $\mathcal{Z}_{p}$ is discrete (dimension 0) for each $p>1$.
@@
By Sternberg–Williams–Ziemer [51, Theorem 3.4], least gradient functions in dimension 3 have critical sets of Hausdorff dimension at most 1 (consisting of isolated points and possibly curves connecting boundary components).
@@
    Uniform bound on cardinality via index theory: The index theory for $p$-harmonic functions developed by Aronsson–Lindqvist [9, Theorem 5.1] provides a topological bound on the number of critical points. For a $p$-harmonic function $u:M\to[0,1]$ with Dirichlet boundary conditions, the Poincaré–Hopf theorem applied to the gradient vector field $\nabla u$ yields:
@@
    where $\chi(M,\partial M)$ is the Euler characteristic of the manifold with boundary. For our geometry $\tilde{M}\cong[0,1]\times S^{2}$ with $\partial\tilde{M}=\{0\}\times S^{2}$, we have $\chi(\tilde{M},\partial\tilde{M})=\chi(S^{2})=2$. Since critical points of capacitary potentials are saddle points with index $\pm 1$ [37, Proposition 4.3], this bounds $|\mathcal{Z}_{p}|\leq 2$ independent of $p$. More generally, $|\mathcal{Z}_{p}|\leq C(\chi(M))$ where $C$ depends only on the topology of $M$.
@@
    Limit of critical points: By uniform $C^{1,\alpha}$ bounds (Lemma 6.28), a subsequence $u_{p_{k}}\to u_{1}$ in $C^{1}$. If $x_{k}\in\mathcal{Z}_{p_{k}}$ with $x_{k}\to x_{*}$, then $\nabla u_{1}(x_{*})=\lim_{k}\nabla u_{p_{k}}(x_{k})=0$, so $x_{*}\in\mathcal{Z}_{1}$.
@@
$\|u_{p,\epsilon}\|_{C^{1,\alpha}(K)}\leq C(K)\quad\text{uniformly in }p\in(1,2],\,\epsilon\in(0,1],$
@@
The crucial point is that this stability constant $C$ depends on the $C^{1,\alpha}$ norm of $u_{p}$, which is uniformly bounded in $p\in(1,2]$ by Lemma 6.28 and Lemma 6.30. Specifically:
  • Lemma 6.28 provides $\|u_{p}\|_{C^{1,\alpha}(K)}\leq C(K)$ uniformly in $p$;
@@
where the uniformity in $p$ follows from the uniform bounds on $\|u_{p}\|_{C^{1,\alpha}}$, $A_{p}(t)$, and $\text{Perimeter}(\Sigma_{t})$.
@@
    Tolksdorf $C^{1,\alpha}$ constant: From [52, Theorem 1.1], for $p$-harmonic $u$ on a domain $\Omega$ with $|\nabla u|\geq c_{0}>0$, the Hölder constant satisfies
    $u]_{C^{1,\alpha}(K)}\leq C_{T}(n,c_{0}/\|\nabla u\|_{\infty})\cdot\|\nabla u\|_{L^{\infty}(\Omega)}$
@@
    for some $\gamma>0$. The exponent $\gamma$ arises from the Hölder interpolation in the Arzelà–Ascoli compactness argument: specifically, $\gamma=\alpha^{\prime}/(\alpha^{\prime}+1)$ where $\alpha^{\prime}<\alpha$ is the limiting Hölder exponent from (C1). This explicit rate ensures quantitative convergence in the double limit, not merely existence.
@@
    Critical set dimension (uniform in $p$): By [80, Theorem 1.2] (extending [81]), the critical set $\mathcal{C}_{p}=\{|\nabla u_{p}|=0\}$ satisfies
@@
    Crucially, the Hausdorff dimension bound depends only on the ellipticity ratio and domain geometry, not on $p$. This ensures the measure of level sets intersecting $\mathcal{C}_{p}$ remains negligible uniformly in $p$.
@@
    Explicit Moore–Osgood verification: For the double limit $\lim_{p\to 1^{+}}\lim_{\epsilon\to 0^{+}}A_{p,\epsilon}(t)=\lim_{\epsilon\to 0^{+}}\lim_{p\to 1^{+}}A_{p,\epsilon}(t)$, we verify Moore–Osgood hypotheses:
@@
Theorem 6.33 (Global AM-Hawking Mass Bound).
Under the hypotheses of Theorem 1.2, the AM-Hawking mass functional is monotonically non-decreasing in $t$:
$m_{H,J}(t_{1})\leq m_{H,J}(t_{2})\quad\text{for all }0\leq t_{1}\leq t_{2}\leq 1,$
and satisfies the global bound:

### v3

@@ 6.4 Monotonicity
Conclusion. Combining Steps 5–7, we obtain uniform $C^{1,\beta}$ bounds for some $\alpha>0$ (possibly small but positive), independent of $p\in(1,2]$. ∎
Remark 6.40 (Explicit Quantitative Bounds for the $p\to 1^{+}$ Limit).
We summarize the key quantitative estimates used in the $p\to 1^{+}$ limit, with explicit dependence on parameters:
  1. 1.
[…]
    $\|\nabla u_{p}\|_{L^{\infty}(\tilde{M})}\leq C_{1}(\tilde{g},\Sigma)\quad\text{uniformly in }p\in(1,2].$
    This follows from the comparison principle: $|\nabla u_{p}|\leq\|\nabla G\|_{L^{\infty}}$ where $G$ is the Green’s function-like comparison function.
  2. 2.
    Gradient lower bound away from critical set: For any $\delta>0$:
    $|\nabla u_{p}(x)|\geq c_{0}(\delta,\tilde{g})>0\quad\text{for }\mathrm{dist}(x,\mathcal{Z}_{p})\geq\delta,\text{ uniformly in }p\in(1,2].$
[…]
  3. 3.
    Hölder exponent: The Lieberman Hölder exponent satisfies:
    $\alpha(p)\geq\alpha_{0}\cdot\min\left(1,(p-1)\left(\frac{c_{0}}{C_{1}}\right)^{2-p}\right),$
    where $\alpha_{0}\in(0,1)$ is the limiting ($p=2$) Hölder exponent. For $p$ close to 1:
    $\alpha(p)\geq\alpha_{0}(p-1)\quad\text{(linear in }p-1).$
    Although $\alpha(p)\to 0$ as $p\to 1^{+}$, the uniform $C^{1,0}$ (Lipschitz) bounds suffice for compactness.
  4. 4.
    Compactness: The family $\{u_{p}\}_{p\in(1,2]}$ is precompact in $C^{1}(K)$ for any compact $K\subset\tilde{M}\setminus\mathcal{Z}$ by Arzelà–Ascoli applied to $\nabla u_{p}$:
    • Uniform boundedness: $\|\nabla u_{p}\|_{L^{\infty}(K)}\leq C_{1}$;
    • Equicontinuity: $|\nabla u_{p}(x)-\nabla u_{p}(y)|\leq C_{2}c_{0}^{-1}C_{1}^{2}|x-y|$ (DiBenedetto estimate).
  5. 5.
    Monotonicity constant: The monotonicity formula coefficient $\frac{d}{dt}m_{H,J}^{2}(t)\geq 0$ holds with explicit lower bound:
    $\frac{d}{dt}m_{H,J}^{2}(t)\geq\frac{1}{16\pi}\int_{\Sigma_{t}}\left(R_{\tilde{g}}-|H|^{2}-W\right)|\nabla u_{p}|^{-1}dA,$
    where $R_{\tilde{g}}\geq 0$ by construction. The bound is independent of $p$ (given uniform control on $|\nabla u_{p}|^{-1}$).
Verification checklist: The reader can verify these bounds by:
  • Gradient $L^{\infty}$: Tolksdorf [56, Theorem 2.1], comparison with barrier functions;
  • Gradient lower bound: Harnack inequality [54, Theorem 1.2] + connectivity;
  • Hölder exponent: Lieberman [34, Section 2], tracking constants in the proof;
  • Equicontinuity: DiBenedetto [24, Chapter VIII, Theorem 1.1].
@@
    $C^{1,\beta}$ regularity: $\|u_{p}\|_{C^{1,\beta}(K)}\leq C(K)$ uniformly in $p\in(1,2]$ (Lemma 6.39);
@@
These three bounds ensure that the Tolksdorf stability estimate for $p$-harmonic functions [56, Theorem 3.2] applies with constants independent of $p$, validating the Moore–Osgood double limit interchange in Remark 6.37.
@@
(i) By the Heinonen–Kilpeläinen–Martio structure theorem [30, Theorem 7.46], the critical set $\mathcal{Z}_{p}$ of a $p$-harmonic function in dimension 3 has Hausdorff dimension at most 1. For the AMO capacitary potential with Dirichlet boundary conditions, the classification of singularities (Manfredi [36]) shows that critical points are saddle points, which are isolated for capacitary potentials. Therefore $\mathcal{Z}_{p}$ is discrete (hence has measure zero). The set of critical values $\{t:\exists x\in u_{p}^{-1}(t)\text{ with }\nabla u_{p}(x)=0\}$ is at most countable, thus has measure zero in $[0,1]$.
Note: The classical Sard theorem requires $C^{n}$ regularity for functions on $n$-dimensional manifolds, which $p$-harmonic functions (being only $C^{1,\beta}$) do not satisfy. The above argument uses the specialized structure theory for $p$-harmonic equations instead.
(ii) Away from $\mathcal{Z}_{p}$, the $p$-harmonic equation is uniformly elliptic. The Harnack inequality for $p$-harmonic functions [54, Theorem 1.2] gives:
@@
AMO boundary conditions exclude critical curves. For the AMO potential $u_{p}:\tilde{M}\to[0,1]$ with $u_{p}|_{\Sigma}=0$ and $u_{p}\to 1$ at infinity, we have stronger control. The key observation is that $u_{p}$ is a capacitary potential—it minimizes the $p$-energy among functions with the given boundary values. By Manfredi [36, Theorem 4.1], capacitary potentials in dimension 3 have critical sets of dimension $\leq 0$ (isolated points) when the boundary data is “generic” in the sense that no boundary component has vanishing $p$-capacity.
@@
Saddle points of capacitary potentials are isolated by the classification of singularities in Aronsson–Lindqvist [8, Section 5]. Therefore $\mathcal{Z}_{p}$ is discrete (dimension 0) for each $p>1$.
@@
By Sternberg–Williams–Ziemer [55, Theorem 3.4], least gradient functions in dimension 3 have critical sets of Hausdorff dimension at most 1 (consisting of isolated points and possibly curves connecting boundary components).
@@
    Uniform bound on cardinality via index theory: The index theory for $p$-harmonic functions developed by Aronsson–Lindqvist [8, Theorem 5.1] provides a topological bound on the number of critical points. For a $p$-harmonic function $u:M\to[0,1]$ with Dirichlet boundary conditions, the Poincaré–Hopf theorem applied to the gradient vector field $\nabla u$ yields:
@@
    where $\chi(M,\partial M)$ is the Euler characteristic of the manifold with boundary. For our geometry $\tilde{M}\cong[0,1]\times S^{2}$ with $\partial\tilde{M}=\{0\}\times S^{2}$, we have $\chi(\tilde{M},\partial\tilde{M})=\chi(S^{2})=2$. Since critical points of capacitary potentials are saddle points with index $\pm 1$ [36, Proposition 4.3], this bounds $|\mathcal{Z}_{p}|\leq 2$ independent of $p$. More generally, $|\mathcal{Z}_{p}|\leq C(\chi(M))$ where $C$ depends only on the topology of $M$.
@@
    Limit of critical points: By uniform $C^{1,\beta}$ bounds (Lemma 6.39), a subsequence $u_{p_{k}}\to u_{1}$ in $C^{1}$. If $x_{k}\in\mathcal{Z}_{p_{k}}$ with $x_{k}\to x_{*}$, then $\nabla u_{1}(x_{*})=\lim_{k}\nabla u_{p_{k}}(x_{k})=0$, so $x_{*}\in\mathcal{Z}_{1}$.
@@
Remark 6.44 (Critical Clarification: “For a.e. tt” vs. “For all tt”).
We clarify which parts of the monotonicity hold for a.e. $t$ versus for all $t$, and why this is sufficient.
(1) What holds for a.e. $t$:
  • The level sets $\Sigma_{t}=\{u_{p}=t\}$ are smooth embedded surfaces for a.e. $t\in(0,1)$ (by the critical set structure theory for $p$-harmonic functions, Remark 6.3).
  • The derivative formula $\frac{d}{dt}m_{H,J}^{2}(t)\geq 0$ holds for a.e. $t$ (at regular values where $\nabla u_{p}\neq 0$ on $\Sigma_{t}$).
  • The area and Willmore functionals $A(t)$, $W(t)$ are differentiable for a.e. $t$.
(2) What holds for ALL $t$:
  • The functions $t\mapsto A(t)$, $t\mapsto m_{H}(t)$, $t\mapsto m_{H,J}(t)$ are continuous and absolutely continuous on $[0,1]$.
  • The boundary values $m_{H,J}(0)$ and $m_{H,J}(1)$ are well-defined as limits.
  • The monotonicity $m_{H,J}(t_{1})\leq m_{H,J}(t_{2})$ for $t_{1}<t_{2}$ holds for ALL $t_{1},t_{2}\in[0,1]$ (including critical values).
(3) Why a.e. suffices for the inequality: The key is the fundamental theorem of calculus for absolutely continuous functions. Since $m_{H,J}^{2}(t)$ is absolutely continuous and $\frac{d}{dt}m_{H,J}^{2}\geq 0$ for a.e. $t$:
$m_{H,J}^{2}(1)-m_{H,J}^{2}(0)=\int_{0}^{1}\frac{d}{dt}m_{H,J}^{2}(t)\,dt\geq 0.$
The singular set $\{t:\nabla u_{p}=0\text{ somewhere on }\Sigma_{t}\}$ has measure zero (by the Heinonen–Kilpeläinen–Martio structure theorem [30]), so its contribution to the integral vanishes. Therefore:
$m_{H,J}(1)\geq m_{H,J}(0)\quad\text{holds unconditionally}.$
(4) Why critical points do not obstruct: At a critical value $t_{*}$ where $\Sigma_{t_{*}}$ contains a critical point, the level set may have singularities (non-smooth points). However:
  • By Lemma 6.42(iv), critical points are isolated (dimension 0).
  • The area $A(t_{*})$ and Hawking mass $m_{H}(t_{*})$ remain finite (the singularity is removable for these integral quantities).
  • The one-sided limits $\lim_{t\to t_{*}^{\pm}}m_{H,J}(t)$ exist and agree, establishing continuity through critical values.
Conclusion: The “a.e. tt” condition is technically necessary for the pointwise derivative formula, but global monotonicity $m_{H,J}(1)\geq m_{H,J}(0)$ holds unconditionally by integration.
Remark 6.45 (Regularity at Critical Points—Detailed Analysis).
@@
$\|u_{p,\epsilon}\|_{C^{1,\beta}(K)}\leq C(K)\quad\text{uniformly in }p\in(1,2],\,\epsilon\in(0,1],$
@@
The essential point is that this stability constant $C$ depends on the $C^{1,\beta}$ norm of $u_{p}$, which is uniformly bounded in $p\in(1,2]$ by Lemma 6.39 and Lemma 6.42. Specifically:
  • Lemma 6.39 provides $\|u_{p}\|_{C^{1,\beta}(K)}\leq C(K)$ uniformly in $p$;
@@
where the uniformity in $p$ follows from the uniform bounds on $\|u_{p}\|_{C^{1,\beta}}$, $A_{p}(t)$, and $\text{Perimeter}(\Sigma_{t})$.
@@
    Tolksdorf $C^{1,\beta}$ constant: From [56, Theorem 1.1], for $p$-harmonic $u$ on a domain $\Omega$ with $|\nabla u|\geq c_{0}>0$, the Hölder constant satisfies
    $u]_{C^{1,\beta}(K)}\leq C_{T}(n,c_{0}/\|\nabla u\|_{\infty})\cdot\|\nabla u\|_{L^{\infty}(\Omega)}$
@@
    for some $\gamma>0$ depending on the Arzelà–Ascoli extraction, which ensures finite iteration of the double limit.
@@
    Critical set dimension (uniform in $p$): By [91, Theorem 1.2] (extending [92]), the critical set $\mathcal{C}_{p}=\{|\nabla u_{p}|=0\}$ satisfies
@@
    The Hausdorff dimension bound depends only on the ellipticity ratio and domain geometry, not on the specific value of $p$. This ensures the measure of level sets intersecting $\mathcal{C}_{p}$ remains negligible uniformly in $p$.
@@
    Explicit Moore–Osgood verification: For the double limit $\lim_{p\to 1^{+}}\lim_{\epsilon\to 0^{+}}A_{p,\epsilon}(t)=\lim_{\epsilon\to 0^{+}}\lim_{p\to 1^{+}}A_{p,\epsilon}(t)$, we verify Moore–Osgood hypotheses explicitly:
@@
Theorem 6.47 (Limit Passage $p\to 1^{+}$: Consolidated Statement).
[…]
Part A: Uniform Bounds (independent of $p$).
  1. (U1)
[…]
  2. (U2)
    Gradient lower bound away from critical set: For any $\delta>0$, there exists $c_{0}(\delta)>0$ such that
    $|\nabla u_{p}(x)|\geq c_{0}(\delta)\quad\text{whenever }\mathrm{dist}(x,\mathcal{Z}_{p})\geq\delta,$
    where $\mathcal{Z}_{p}:=\{x\in\tilde{M}:\nabla u_{p}(x)=0\}$ and $c_{0}(\delta)$ is independent of $p$.
  3. (U3)
    Hölder regularity: For any compact $K\subset\tilde{M}$ with $\mathrm{dist}(K,\partial\tilde{M})>0$:
    $\|u_{p}\|_{C^{1,\beta}(K)}\leq C_{2}(K)\quad\text{for all }p\in(1,2],$
    where $\alpha=\alpha(n,c_{0}/C_{1})\in(0,1)$ and $C_{2}(K)$ are independent of $p$ (Lemma 6.39).
  4. (U4)
    Critical set structure: $\dim_{\mathcal{H}}(\mathcal{Z}_{p})\leq 1$ for all $p\in(1,2]$, and $|\mathcal{Z}_{p}|\leq N_{\mathrm{top}}$ where $N_{\mathrm{top}}$ depends only on the topology of $\tilde{M}$ (Lemma 6.42(iv)).
  5. (U5)
    Area and mass bounds: For a.e. $t\in(0,1)$:
    $A_{0}\leq A_{p}(t)\leq C_{3},\quad|m_{H,J,p}(t)|\leq C_{4},$
    where $A_{0}>0$ is the horizon area and $C_{3},C_{4}$ depend only on the initial data.
Part B: Convergence Mode.
  1. (C1)
    $C^{1,\alpha^{\prime}}$ locally uniform convergence: For any $\alpha^{\prime}<\alpha$ and compact $K\subset\tilde{M}\setminus\mathcal{Z}_{1}$:
[…]
  2. (C2)
[…]
  3. (C3)
    Level set convergence: For a.e. $t\in(0,1)$, the level sets $\Sigma_{t}^{(p)}:=\{u_{p}=t\}$ converge to $\Sigma_{t}^{(1)}:=\{u_{1}=t\}$ in the Hausdorff metric.
  4. (C4)
    Functional convergence: $A_{p}(t)\to A_{1}(t)$ and $m_{H,J,p}(t)\to m_{H,J,1}(t)$ uniformly on compact subsets of $(0,1)\setminus T_{\mathrm{crit}}$, where $T_{\mathrm{crit}}:=\{t:t\in u_{1}(\mathcal{Z}_{1})\}$ has measure zero.
Part C: Passage of “a.e. in tt” Statements to the Limit.
  1. (L1)
    Preservation of monotonicity: The derivative inequality
    $\frac{d}{dt}m_{H,J,p}^{2}(t)\geq 0\quad\text{for a.e.\ }t\in(0,1)$
    holds for each $p\in(1,2]$. Taking $p\to 1^{+}$:
    $\frac{d}{dt}m_{H,J,1}^{2}(t)\geq 0\quad\text{for a.e.\ }t\in(0,1)$
    in the distributional sense.
  2. (L2)
    Global monotonicity via absolute continuity: Since $t\mapsto m_{H,J,p}(t)$ is absolutely continuous for each $p$ (by the co-area formula), and absolute continuity is preserved under locally uniform limits, the function $t\mapsto m_{H,J,1}(t)$ is absolutely continuous. Combined with (L1):
    $m_{H,J,1}(t_{2})\geq m_{H,J,1}(t_{1})\quad\text{for ALL }0\leq t_{1}<t_{2}\leq 1.$
  3. (L3)
    Exceptional set control: The set of $t$ where the pointwise derivative formula fails satisfies
    $\mathcal{E}:=\{t\in(0,1):\Sigma_{t}^{(1)}\text{ is singular}\}\subset T_{\mathrm{crit}},$
    which has measure zero uniformly in the approximation. Thus the “a.e.” condition is not weakened in the limit.
Proof.
Part A: (U1) follows from the maximum principle for $p$-harmonic functions with bounded boundary data. (U2)–(U3) are Lemma 6.42(ii) and Lemma 6.39 respectively, whose proofs establish uniformity via the Tolksdorf and Lieberman estimates. (U4) combines Heinonen–Kilpeläinen–Martio [30, Theorem 7.46] with the index bound from Lemma 6.42(iv). (U5) follows from the isoperimetric inequality and the co-area formula.
Part B: (C1) is Arzelà–Ascoli applied to the equicontinuous family $\{u_{p}\}_{p\in(1,2]}$ with (U3). (C2) follows from the energy bound $\int|\nabla u_{p}|\,dV\leq C$ and weak compactness. (C3) is a consequence of (C1) at regular values. (C4) follows from (C3) and the continuity of area/mass functionals.
Part C: (L1) follows from the weak convergence of non-negative measures: if $\mu_{p}:=(d/dt)m_{H,J,p}^{2}\cdot\mathcal{L}^{1}\geq 0$ as measures, then any weak-* limit $\mu_{1}$ satisfies $\mu_{1}\geq 0$. (L2) is the fundamental theorem of calculus for absolutely continuous functions. (L3) uses the uniform bound (U4) and the fact that $T_{\mathrm{crit}}=u_{1}(\mathcal{Z}_{1})$ has measure zero by Sard’s theorem applied to the Lipschitz function $u_{1}|_{\mathcal{Z}_{1}}$. ∎
Remark 6.48 (Why This Consolidation Matters).
The limit $p\to 1^{+}$ is the technical heart of the AMO approach. The estimates scattered throughout this section (Lemma 6.39, Lemma 6.42, Remarks 6.446.46) are now collected in Theorem 6.47 to make explicit:
  1. 1.
    Which quantities are uniformly bounded (Part A) — essential for compactness arguments;
  2. 2.
    In what topology convergence occurs (Part B) — $C^{1,\alpha^{\prime}}$ locally, not merely $L^{p}$;
  3. 3.
    How “a.e. in tt” passes to the limit (Part C) — via absolute continuity, not pointwise limits of exceptional sets.
The key insight is that the “a.e.” condition does not degrade under limits: the exceptional set $T_{\mathrm{crit}}$ remains measure-zero because critical points cannot accumulate (Part A, (U4)), and absolute continuity converts the a.e. derivative bound into global monotonicity.
Theorem 6.49 (Rigorous AM-Hawking Monotonicity).
Under the hypotheses of Theorem 1.2, the AM-Hawking mass functional satisfies: